# Testing Seesaw and Leptogenesis via Gravitational Waves: Majorana versus Dirac

### v1

@@ Abstract
We investigate the B-L gauge extension of the Standard Model that the Dirac seesaw mechanism with thermal Leptogenesis can be tested using the stochastic gravitational background (SGWB) emanating from a network of cosmic strings when B-L symmetry is broken. With right-handed neutrino mass lighter than the typical scale of grand unification, the B-L symmetry protecting the right-handed neutrinos leads to constraints on the Yukawa couplings for both Dirac and Majorana scenarios. Estimating the predicted gravitational wave background we find that future space-borne missions could probe the range concerning thermal Dirac Leptogenesis. In a comparative analysis between such probes of gravitational wave sourced from cosmic strings in Dirac and Majorana Leptogenesis in the B-L extension, based on the energy scales of the Leptogenesis, for instance, GW detectors will be able to probe the scale of Dirac Leptogenesis upto $10^{9}$ GeV, while for Majorana Leptogenesis it would be upto $10^{12}$ GeV.
@@ I Introduction
The evidence of going beyond the Standard Model (SM) of particle physics arises undisputedly from neutrino oscillation experiments involving solar Super-Kamiokande:2001bfk ; Super-Kamiokande:2002ujc ; SNO:2002tuh ; Super-Kamiokande:2005mbp ; Super-Kamiokande:2016yck ; Borexino:2015axw , atmospheric IceCube:2017lak ; ANTARES:2018rtf and reactor KamLAND:2008dgz ; T2K:2011ypd ; DoubleChooz:2011ymz ; T2K:2013ppw . We infer from these experiments that neutrinos have mass and the flavor states mix due to the propagation of multiple mass eigenstates. This however only quantifies the squared mass differences of the neutrinos and not the absolute mass scales involved. On the other hand, the $\beta$-decay experiment KATRIN KATRIN:2021uub gives a stringent direct limit on the absolute value of neutrino mass scale, $m_{\nu}<0.8$ eV.
On the cosmological frontiers, measurements of the Cosmic Microwave Background radiation (CMBR) by Planck 2018 Planck:2018vyg and large-scale structure (LSS) constrain the sum of all neutrino masses to $\sum_{i}m_{\nu_{i}}<0.12$ eV Planck:2018vyg ; eBOSS:2020yzd . On the other hand, the observed baryon asymmetry of the Universe (BAU) Zyla:2020zbs ; Planck:2018vyg , often expressed in terms of the baryon to photon ratio is, according to the Planck data Planck:2018vyg  [^1] It is also expressed as $Y_{B}=\frac{n_{\bar{B}}}{s}=(0.82-0.92)\times 10^{-10}$. We use the central value $Y_{B}=\frac{n_{\bar{B}}}{s}=8.66\times 10^{-11}$ in our estimates.,
@@
This agrees with the values of abundances of light elements extracted from BBN (Big Bang Nucleosynthesis) data Fields:2019pfx .
@@
Just augmenting the Standard Model with two or more right-handed Majorana neutrinos (RHNs), one can explain the tiny neutrino masses, generated via the well-known Type-I seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Glashow:1979nm ; Mohapatra:1979ia , while baryon asymmetry can be explained using the baryogenesis via Leptogenesis mechanism Fukugita:1986hr , provides a unified explanation. Nonetheless this mechanism demands the lightest RHN mass to exceed the Davidson-Ibarra bound, $M_{1}\gtrsim 10^{9}$ GeV, assuming the RHN mass spectrum is hierarchical, along with no lepton flavor effects Nardi:2006fx ; Abada:2006fw ; Abada:2006ea ; Davidson:2002qv ; Giudice:2003jh . This deflates any hope of direct detection of the high-scale seesaw and Leptogenesis since they lie well beyond the energy reaches of the current and future laboratory experiments.
@@
Nonetheless, certain indirect signatures of new physics like lepton number violation processes through neutrinoless double beta decay Cirigliano:2022oqy or CP violation in neutrino oscillation Endoh:2002wm are searched for in laboratories. There can be certain theoretical constraints on the low energy values of couplings, whose consistency conditions for UV-completion such as in $SO(10)$ Grand Unified Theories (GUT)  DiBari:2008mp ; Bertuzzo:2010et ; Buccella:2012kc ; Altarelli:2013aqa ; Fong:2014gea ; Mummidi:2021anm ; Patel:2022xxu or with the demand of electroweak (EW) vacuum (meta)stability in the early Universe Ipek:2018sai ; Croon:2019dfw sometimes provide some useful bounds on the large parameter space involved in Seesaw and Leptogenesis scenarios.
@@
The discovery of Gravitational Waves (GWs) from black hole mergers by LIGO, Virgo, KAGRA collaboration LIGOScientific:2016aoc ; LIGOScientific:2016sjg and evidence for measurements of stochastic GW background by pulsar timing array (PTA) observations Carilli:2004nx ; Janssen:2014dka ; Weltman:2018zrl ; EPTA:2015qep ; EPTA:2015gke ; NANOGrav:2023gor ; NANOGrav:2023hvm have lead to several new physics cases to detect GWs of primordial and cosmological origins as well. In the context of baryogenesis via Leptogenesis scenarios investigated so far, cosmological pathways to probe such high-scale physics involve predictions of the CMB spectral indices Ghoshal:2022fud or gravitational waves from local cosmic strings Dror:2019syi ; Saad:2022mzu ; DiBari:2023mwu ; Blasi:2020wpy , global cosmic strings Fu:2023nrn , domain walls Barman:2022yos ; King:2023cgv , nucleating and colliding vacuum bubbles Dasgupta:2022isg ; Borah:2022cdx ; Ghosh:2025non , graviton bremmstrahlung Ghoshal:2022kqp , inflationary tensor perturbations of first-order Berbig:2023yyy ; Borboruah:2024eal and second-order Afzal:2024xci ; Chianese:2025mll , primordial blackholes Perez-Gonzalez:2020vnz ; Datta:2020bht ; JyotiDas:2021shi ; Barman:2021ost ; Bernal:2022pue ; Bhaumik:2022pil ; Bhaumik:2022zdd ; Ghoshal:2023fno , and utisilizing features of the primordial non-gaussianity and curvature bi-spectrum and tri-spectrum in CMB Cui:2021iie ; Fong:2023egk . Particularly, it has been shown that one may probe directly as a possible foorptint of high-scale leptogenesis with GW from cosmic strings, via investigating only the RHN mass sector, the origin of which may occur at a more fundamental level due to symmetry breaking Dror:2019syi ; Blasi:2020wpy ; Samanta:2020cdk ; Datta:2020bht ; Chianese:2024gee . For instance, the RHNs become massive because of the spontaneous breaking of a $U(1)_{B-L}$ gauge invariance, a symmetry which quite naturally gets embedded in several Grand Unified Theories (GUT) theories Davidson:1978pm ; Marshak:1979fm ; Buchmuller:2013lra ; Buchmuller:2019gfy , and a breaking of which inevitably produces cosmic string network that radiates gravitational wave signals. For any $U(1)_{B-L}$ invariant RHN mass term, the Yukawa coupling $y_{N}\overline{N_{R}}\Phi N^{C}_{R}$, where $y_{N}$ is the Yukawa coupling, $N_{R}$ represents the RHN particle and $\Phi$ denotes the BLB-L scalar with vacuum expectation value $v_{\Phi}$. When the $U(1)_{B-L}$ symmetry is broken, the RHNs obtain mass given by $M_{N}=y_{N}v_{\Phi}$, thereby relating to the properties (like string tension) of cosmic strings which are formed after the symmetry breaking Kibble:1976sj ; Hindmarsh:1994re ; Jeannerot:2003qv leading to production of GWs with a almost scale-invariant spectrum with an overall GW amplitude $\Omega_{\rm GW}\propto v_{\Phi}$ Vilenkin:1981bx ; Vachaspati:1984gt .
@@
Under this circumstance, while most of the Leptogenesis scenarios consider Majorana nature of light neutrinos, one equally appealing alternative is to consider Dirac nature of light neutrinos. As originally proposed in Refs. Dick:1999je ; Murayama:2002je , one can have successful Leptogenesis even with light Dirac neutrino scenarios where total lepton number or BLB-L is conserved just like it is conversed in the SM of particle physics  [^2] The recently proposed mechanism of wash-in Leptogenesis Domcke:2020quw ; Domcke:2022kfs advocates an approach getting rid of CP violation in the RHN sector altogether and separates the energy scales of CP violation and B-L violation altogether in presence of non-trivial chemical background configurations. Like Dirac Leptogenesis, here the RHNs mere spectator processes that reprocess the chemical potentials of the SM particle species in the thermal bath. The action of the electroweak sphalerons on the chemical composition of the SM plasma then results in the usual violation of baryon-plus-lepton number B+L, while the action of the RHN interactions results in the violation of B-L. . This mechanism is popularly known as the Dirac Leptogenesis. What this mechanism exploits is the creation of an equal and opposite amount of lepton asymmetry in left handed and right handed neutrino sectors which is subsequently followed by the conversion of left sector asymmetry into baryon asymmetry via the electroweak sphalerons. This is made possible since the right-handed sector is invisible to the sphalerons inside the SM plasma, only the left-handed asymmetry is converted into baryons. The asymmetry within the $\nu_{R}$ sector, and indeed the $\nu_{R}$ particles themselves, are well-high impossible to observe due to their extreme weak interactions. Needless to mention here, the effective Dirac Yukawa couplings are naturally very tiny which guarantee that the lepton asymmetries in the left and right handed sectors do not equibriliate and washed out. A myriad of possible implementation of this idea have been considered, see some in Refs. Boz:2004ga ; Thomas:2005rs ; Thomas:2006gr ; Cerdeno:2006ha ; Gu:2006dc ; Gu:2007mc ; Chun:2008pg ; Bechinger:2009qk ; Chen:2011sb ; Choi:2012ba ; Borah:2016zbd ; Gu:2016hxh ; Narendra:2017uxl . Particularly, in a few related works Heeck:2013vha ; Gu:2019yvw ; Mahanta:2021plx , violation of BLB-L symmetry was accommodated in a way to preserve the Dirac nature of light neutrinos, while generating lepton asymmetry simultaneously.
@@
In this paper we first discuss Dirac Leptogenesis, with thermalized RHNs and therefore suitable within the neutrinophilic Ma:2000cc two-Higgs-doublet which naturally explains the smallness of Dirac masses Wang:2006jy ; Gabriel:2006ns ; Davidson:2009ha ; Davidson:2010sf . Moreover we follow the framework of lepton-number-violating (LNV) Dirac neutrinos Heeck:2013rpa ; Heeck:2013vha to create a lepton asymmetry from the $CP$-violating decay of a heavy particle that is present in the seesaw model. [^3] Prior to Ref. Heeck:2013rpa , Ref. Chen:2012jg mentioned roles of LNV Dirac neutrinos in the early Universe. Due to this reason, the mechanism is actually more close to standard Leptogenesis than neutrinogenesis, despite the fact that it contains Dirac neutrinos. Finally we do a detailed investigation of the parameter space of leptogenesis, in B-L extensions of the SM, comparing Majorana and Dirac leptogenesis. In terms of scales of seesaw and Leptogensis, we provide quantitative estimates regarding various GW detectors like LISA and ET which are able to measure GW from B-L symmetry breaking cosmic strings, probing large regions of such parameter space in consists of different correlations between seesaw scale and the scales of Dirac and Majorana leptogenesis.
@@ II The B-L extended Model
We extend the SM symmetry with a well motivated beyond the SM framework based on the gauged $U(1)_{B-L}$ symmetry, where $B$ and $L$ are the baryon and lepton numbers respectively. Considering the gauged $U(1)_{B-L}$ symmetry brings unwanted triangular anomalies to the framework. These anomalies can be canceled by introducing three right-handed neutrinos to the framework, which is also essential for realizing the neutrino masses. The overall particle content of our model extend with three right-handed neutrinos $\nu_{R_{\alpha}},\,\alpha=\{1,2,3\}$, two singlet scalars $\varphi$ and $\xi$, under the SM gauge group. The scalar $\varphi$ is introduced in order to spontaneously break the BLB-L symmetry by four units. The $\xi$ acts as a mediator which is responsible for the creation of asymmetry in the right-handed neutrinos Heeck:2013vha . The BLB-L charge assignment of the new particle content can be seen in Table 1 .
@@
The $\mathcal{L}_{\mathrm{SM}},\,\mathcal{L}_{\mathrm{kinetic}},\,\mathcal{L}_{Z^{\prime}}$ are the all SM particle interactions, kinetic terms of new particle content and $Z^{\prime}$ boson interaction terms, respectively. $H$ is the SM Higgs doublet. We consider $\mu_{H}^{2},\,\mu_{\varphi}^{2}<0<\mu_{\xi}^{2}$. Hence the potential has minimum with $\langle\xi\rangle=0$, $\langle H\rangle\neq 0\neq\langle\varphi\rangle$, which breaks the $SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$ to $U(1)_{\text{EM}}\times Z_{4}^{L}$. Under $Z_{4}^{L}$ the leptons transform as $\ell\rightarrow-i\,\ell$ and $\xi\rightarrow-\xi$. The scalar $\xi$ does not acquire a vacuum expectation value (vev) and forbids Majorana mass term while allowing small Dirac neutrino mass $m_{\nu_{\alpha\beta}}=y_{\alpha\beta}\langle H\rangle$ with the Yukawa coupling of the order $\lesssim 10^{-11}$. The Yukawa coupling $\kappa_{\alpha\beta}$ is a complex symmetric matrix, in general. The $\kappa_{\alpha\beta}$ plays a crucial role in generating the $\nu_{R}$ asymmetry. We define $\xi=(\zeta_{1}+i\,\zeta_{2})/\sqrt{2}$, where $\zeta_{1,2}$ real scalars. The $\mu$ term present in the potential induces a mass-splitting between the real scalars $\zeta_{1,2}$:

### v2

# Testing Seesaw and Leptogenesis Gravitational Waves: Majorana versus Dirac cases [^1] This paper is the original version of the article accepted for publication Phys.Rev.D with the title “Imprints of seesaw and leptogenesis on cosmic-string contributions to the stochastic gravitational wave background: Majorana versus Dirac cases”

@@ Abstract
We investigate the B-L gauge extension of the Standard Model that the Dirac seesaw mechanism with thermal Leptogenesis leave imprints in early universe via the stochastic gravitational background (SGWB) emanating from a network of cosmic strings when B-L symmetry is broken. With right-handed neutrino mass lighter than the typical scale of grand unification, the B-L symmetry protecting the right-handed neutrinos leads to constraints on the Yukawa couplings for both Dirac and Majorana scenarios. Estimating the predicted gravitational wave background we find that future space-borne missions could probe the range concerning thermal Dirac Leptogenesis. In a comparative analysis between such probes of gravitational wave sourced from cosmic strings in Dirac and Majorana Leptogenesis in the B-L extension, based on the energy scales of the Leptogenesis, for instance, GW detectors will be able to probe the scale of Dirac Leptogenesis upto $10^{9}$ GeV, while for Majorana Leptogenesis it would be upto $10^{12}$ GeV.
@@ I Introduction
The evidence of going beyond the Standard Model (SM) of particle physics arises undisputedly from neutrino oscillation experiments involving solar Fukuda and others (2001, 2002); Ahmad and others (2002); Ashie and others (2005); Abe and others (2016); Smirnov and others (2016), atmospheric Aartsen and others (2018); Albert and others (2019) and reactor Abe and others (2008, 2011, 2012, 2014). We infer from these experiments that neutrinos have mass and the flavor states mix due to the propagation of multiple mass eigenstates. This however only quantifies the squared mass differences of the neutrinos and not the absolute mass scales involved. On the other hand, the $\beta$-decay experiment KATRIN Aker and others (2022) gives a stringent direct limit on the absolute value of neutrino mass scale, $m_{\nu}<0.8$ eV.
On the cosmological frontiers, measurements of the Cosmic Microwave Background radiation (CMBR) by Planck 2018 Aghanim and others (2020) and large-scale structure (LSS) constrain the sum of all neutrino masses to $\sum_{i}m_{\nu_{i}}<0.12$ eV Aghanim and others (2020); Alam and others (2021). On the other hand, the observed baryon asymmetry of the Universe (BAU) Zyla and others (2020b); Aghanim and others (2020), often expressed in terms of the baryon to photon ratio is, according to the Planck data Aghanim and others (2020)  [^2] It is also expressed as $Y_{B}=\frac{n_{\bar{B}}}{s}=(0.82-0.92)\times 10^{-10}$. We use the central value $Y_{B}=\frac{n_{\bar{B}}}{s}=8.66\times 10^{-11}$ in our estimates.,
@@
This agrees with the values of abundances of light elements extracted from BBN (Big Bang Nucleosynthesis) data Fields et al. (2020).
@@
Just augmenting the Standard Model with two or more right-handed Majorana neutrinos (RHNs), one can explain the tiny neutrino masses, generated via the well-known Type-I seesaw mechanism Minkowski (1977); Yanagida (1979); Glashow (1980); Mohapatra and Senjanovic (1980), while baryon asymmetry can be explained using the baryogenesis via Leptogenesis mechanism Fukugita and Yanagida (1986), provides a unified explanation. Nonetheless this mechanism demands the lightest RHN mass to exceed the Davidson-Ibarra boundDavidson and Ibarra (2002), $M_{1}\gtrsim 10^{9}$ GeV, assuming the RHN mass spectrum is hierarchical, along with no lepton flavor effects Nardi et al. (2006); Abada et al. (2006b, a); Giudice et al. (2004). This deflates any hope of direct detection of the high-scale seesaw and Leptogenesis since they lie well beyond the energy reaches of the current and future laboratory experiments.
@@
Nonetheless, certain indirect signatures of new physics like lepton number violation processes through neutrinoless double beta decay Cirigliano and others (2022) or CP violation in neutrino oscillation Endoh et al. (2002) are searched for in laboratories. There can be certain theoretical constraints on the low energy values of couplings, whose consistency conditions for UV-completion such as in $SO(10)$ Grand Unified Theories (GUT)  Di Bari and Riotto (2009); Bertuzzo et al. (2011); Buccella et al. (2012); Altarelli and Meloni (2013); Fong et al. (2015); Mummidi and Patel (2021); Patel (2023) or with the demand of electroweak (EW) vacuum (meta)stability in the early Universe Ipek et al. (2018); Croon et al. (2019) sometimes provide some useful bounds on the large parameter space involved in Seesaw and Leptogenesis scenarios.
@@
The discovery of Gravitational Waves (GWs) from black hole mergers by LIGO, Virgo, KAGRA collaboration Abbott and others (2016b, a) and evidence for measurements of stochastic GW background by pulsar timing array (PTA) observations Carilli and Rawlings (2004); Janssen and others (2015); Weltman and others (2020); Lentati and others (2015a); Babak and others (2016b); Agazie and others (2023); Afzal and others (2023) have lead to several new physics cases to detect GWs of primordial and cosmological origins as well. In the context of baryogenesis via Leptogenesis scenarios investigated so far, cosmological pathways to probe such high-scale physics involve predictions of the CMB spectral indices Ghoshal et al. (2024a) or gravitational waves from local cosmic strings Dror et al. (2020); Saad (2023); Di Bari et al. (2024); Blasi et al. (2020), global cosmic strings Fu et al. (2023), domain walls Barman et al. (2022b); King et al. (2024), nucleating and colliding vacuum bubbles Dasgupta et al. (2022); Borah et al. (2022); Ghosh et al. (2025), graviton bremmstrahlung Ghoshal et al. (2023b); Datta and Sil (2024); Datta et al. (2026), inflationary tensor perturbations of first-order Berbig and Ghoshal (2023); Borboruah et al. (2024) and second-order Afzal and Ghoshal (2024); Chianese et al. (2025b), primordial blackholes Perez-Gonzalez and Turner (2021); Datta et al. (2021); Jyoti Das et al. (2021); Barman et al. (2022a); Bernal et al. (2022); Bhaumik et al. (2022, 2023); Ghoshal et al. (2024b), and utisilizing features of the primordial non-gaussianity and curvature bi-spectrum and tri-spectrum in CMB Cui and Xianyu (2022); Fong et al. (2023). Particularly, it has been shown that one may probe directly as a possible foorptint of high-scale leptogenesis with GW from cosmic strings, via investigating only the RHN mass sector, the origin of which may occur at a more fundamental level due to symmetry breaking Dror et al. (2020); Blasi et al. (2020); Samanta and Datta (2021); Datta et al. (2021); Chianese et al. (2025a). For instance, the RHNs become massive because of the spontaneous breaking of a $U(1)_{B-L}$ gauge invariance, a symmetry which quite naturally gets embedded in several Grand Unified Theories (GUT) theories Davidson (1979); Marshak and Mohapatra (1980); Buchmüller et al. (2013); Buchmuller et al. (2020), and a breaking of which inevitably produces cosmic string network that radiates gravitational wave signals. For any $U(1)_{B-L}$ invariant RHN mass term, the Yukawa coupling $y_{N}\overline{N_{R}}\Phi N^{C}_{R}$, where $y_{N}$ is the Yukawa coupling, $N_{R}$ represents the RHN particle and $\Phi$ denotes the BLB-L scalar with vacuum expectation value $v_{\Phi}$. When the $U(1)_{B-L}$ symmetry is broken, the RHNs obtain mass given by $M_{N}=y_{N}v_{\Phi}$, thereby relating to the properties (like string tension) of cosmic strings which are formed after the symmetry breaking Kibble (1976); Hindmarsh and Kibble (1995); Jeannerot et al. (2003) leading to production of GWs with a almost scale-invariant spectrum with an overall GW amplitude $\Omega_{\rm GW}\propto v_{\Phi}$ Vilenkin (1981); Vachaspati and Vilenkin (1985).
@@
Under this circumstance, while most of the Leptogenesis scenarios consider Majorana nature of light neutrinos, one equally appealing alternative is to consider Dirac nature of light neutrinos. As originally proposed in Refs. Dick et al. (2000); Murayama and Pierce (2002), one can have successful Leptogenesis even with light Dirac neutrino scenarios where total lepton number or BLB-L is conserved just like it is conversed in the SM of particle physics  [^3] The recently proposed mechanism of wash-in Leptogenesis Domcke et al. (2021, 2023) advocates an approach getting rid of CP violation in the RHN sector altogether and separates the energy scales of CP violation and B-L violation altogether in presence of non-trivial chemical background configurations. Like Dirac Leptogenesis, here the RHNs mere spectator processes that reprocess the chemical potentials of the SM particle species in the thermal bath. The action of the electroweak sphalerons on the chemical composition of the SM plasma then results in the usual violation of baryon-plus-lepton number B+L, while the action of the RHN interactions results in the violation of B-L, also see Refs.Maleknejad (2014, 2021, 2020) for gravitational production during inflation. . This mechanism is popularly known as the Dirac Leptogenesis. What this mechanism exploits is the creation of an equal and opposite amount of lepton asymmetry in left handed and right handed neutrino sectors which is subsequently followed by the conversion of left sector asymmetry into baryon asymmetry via the electroweak sphalerons. This is made possible since the right-handed sector is invisible to the sphalerons inside the SM plasma, only the left-handed asymmetry is converted into baryons. The asymmetry within the $\nu_{R}$ sector, and indeed the $\nu_{R}$ particles themselves, are well-high impossible to observe due to their extreme weak interactions. Needless to mention here, the effective Dirac Yukawa couplings are naturally very tiny which guarantee that the lepton asymmetries in the left and right handed sectors do not equibriliate and washed out. A myriad of possible implementation of this idea have been considered, see some in Refs. Boz and Pak (2004); Thomas and Toharia (2006, 2007); Cerdeno et al. (2006); Gu and He (2006); Gu et al. (2008); Chun and Roy (2008); Bechinger and Seidl (2010); Chen et al. (2012); Choi et al. (2013); Borah and Dasgupta (2016); Gu (2016); Narendra et al. (2018b). Particularly, in a few related works Heeck (2013); Gu (2020); Mahanta and Borah (2022), violation of BLB-L symmetry was accommodated in a way to preserve the Dirac nature of light neutrinos, while generating lepton asymmetry simultaneously.
@@
In this paper we first discuss Dirac Leptogenesis, with thermalized RHNs and therefore suitable within the neutrinophilic Ma (2001) two-Higgs-doublet which naturally explains the smallness of Dirac masses Wang et al. (2006); Gabriel and Nandi (2007); Davidson and Logan (2009, 2010). Moreover we follow the framework of lepton-number-violating (LNV) Dirac neutrinos Heeck and Rodejohann (2013); Heeck (2013) to create a lepton asymmetry from the $CP$-violating decay of a heavy particle that is present in the seesaw model. [^4] Prior to Ref. Heeck and Rodejohann (2013), Ref. Chen et al. (2013) mentioned roles of LNV Dirac neutrinos in the early Universe. Due to this reason, the mechanism is actually more close to standard Leptogenesis than neutrinogenesis, despite the fact that it contains Dirac neutrinos. Finally we do a detailed investigation of the parameter space of leptogenesis, in B-L extensions of the SM, comparing Majorana and Dirac leptogenesis. In terms of scales of seesaw and Leptogensis, we provide quantitative estimates regarding various GW detectors like LISA and ET which are able to measure GW from B-L symmetry breaking cosmic strings, probing large regions of such parameter space in consists of different correlations between seesaw scale and the scales of Dirac and Majorana leptogenesis. In the current scenario, we do not consider the lepton flavor effects Nardi et al. (2006); Abada et al. (2006b, a), throughout the draft, for simplicity.
@@ II The B-L extended Model
We extend the SM symmetry with a well motivated beyond the SM framework based on the gauged $U(1)_{B-L}$ symmetry, where $B$ and $L$ are the baryon and lepton numbers respectively. Considering the gauged $U(1)_{B-L}$ symmetry brings unwanted triangular anomalies to the framework. These anomalies break gauge symmetry at the quantum level, caused by chiral fermions running in triangle loop diagrams involving gauge bosons. [^5] The triangle anomalies for both $U(1)^{3}_{B-L}$ and the mixed $U(1)_{B-L}-(\text{gravity})^{2}$ diagrams are non-zero. These triangle anomalies for the SM fermion content turns out to be $\mathcal{A}^{\text{SM}}\left[U(1)^{3}_{B-L}\right]=-3$, $\mathcal{A}^{\text{SM}}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]=-3$. These anomalies can be canceled by introducing three right-handed neutrinos to the framework  [^6] The right-handed neutrinos contribute $\mathcal{A}^{\text{New}}\left[U(1)^{3}_{B-L}\right]=3,\mathcal{A}^{\text{New}}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]=3$ leading to vanishing total of triangle anomalies Borah et al. (2020)., which is also essential for realizing the neutrino massesGeng and Marshak (1989); He et al. (1990); Kohri and Sahu (2013); Narendra et al. (2018a); Borah et al. (2020). The overall particle content of our model extend with three right-handed neutrinos $\nu_{R_{\alpha}},\,\alpha=\{1,2,3\}$, two singlet scalars $\varphi$ and $\xi$, under the SM gauge group. The scalar $\varphi$ is introduced in order to spontaneously break the BLB-L symmetry by four units. The $\xi$ acts as a mediator which is responsible for the creation of asymmetry in the right-handed neutrinos Heeck (2013). The BLB-L charge assignment of the new particle content can be seen in Table 1 .
@@
The $\mathcal{L}_{\mathrm{SM}},\,\mathcal{L}_{\mathrm{kinetic}},\,\mathcal{L}_{Z^{\prime}}$ are the all SM particle interactions, kinetic terms of new particle content and $Z^{\prime}$ boson interaction terms, respectively. $H$ is the SM Higgs doublet. We consider $\mu_{H}^{2},\,\mu_{\varphi}^{2}<0<\mu_{\xi}^{2}$. Hence the potential has minimum with $\langle\xi\rangle=0$, $\langle H\rangle\neq 0\neq\langle\varphi\rangle$, which breaks the $SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$ to $U(1)_{\text{EM}}\times Z_{4}^{L}$. Under $Z_{4}^{L}$ the leptons transform as $\ell\rightarrow-i\,\ell$ and $\xi\rightarrow-\xi$. In principle the $\mu$ coupling in the last term in Eq. 3 can induce a non-zero vev for $\xi$ after $\varphi$ acquires a vacuum expectation value (VEV). If $\langle\xi\rangle\neq 0$, it spontaneously breaks the residual symmetry $Z_{4}^{L}$, that would generate Majorana mass terms for $\nu_{R}$, which leads to a standard Majorana Leptogenesis scenario, rather than the intended Dirac Leptogenesis framework. Since $\xi$ does not acquire a VEV and forbids Majorana mass term it allows small Dirac neutrino mass $m_{\nu_{\alpha\beta}}=y_{\alpha\beta}\langle H\rangle$ with the Yukawa coupling of the order $\lesssim 10^{-11}$. The Yukawa coupling $\kappa_{\alpha\beta}$ is a complex symmetric matrix, in general. The $\kappa_{\alpha\beta}$ plays a crucial role in generating the $\nu_{R}$ asymmetry. We define $\xi=(\zeta_{1}+i\,\zeta_{2})/\sqrt{2}$, where $\zeta_{1,2}$ real scalars. The $\mu$ term present in the potential induces a mass-splitting between the real scalars $\zeta_{1,2}$:
@@
The mass scale of $Z_{B-L}$ boson is very heavy, therefore it decouples from the thermal bath early in the Universe, and hence it does not thermalizes the right-handed neutrinos. The gauge interactions mediated by heavy $Z_{B-L}$ boson or the scalar interactions involving $\varphi$ are not sufficient enough to thermalize the RHN.